calculate the molar heat capacity Cm for water. you knew that 30.1 kJ was required to heat 10.0 g of ice from 0.00 C to steam 100.00 C and that the melting point pin ∆fusH = 6.01 kJ / mol and the evaporating enthalpy ∆vapH = 40.7 kJ / mol.
Answer #1

Mass a water = 10 g molas mass of water = 18 g/mol mole of wata = mass mokes mass = 10 If Cm ix Heat capacity per mole Heat B100 lk) 47 - 373(k) – 273(k) - Put all Value 30.1 = (6:01) x0.5555) + (0:5555) Cm (100) + (0.5555)(40.7) 30.1 = 3+3385 + 55.5